# SR-Reward: Taking The Path More Traveled

### v1

@@ Abstract
In this paper, we propose a novel method for learning reward functions directly from offline demonstrations. Unlike traditional inverse reinforcement learning (IRL), our approach decouples the reward function from the learner’s policy, eliminating the adversarial interaction typically required between the two. This results in a more stable and efficient training process. Our reward function, called SR-Reward, leverages successor representation (SR) to encode a state based on expected future states’ visitation under the demonstration policy and transition dynamics. By utilizing the Bellman equation, SR-Reward can be learned concurrently with most reinforcement learning (RL) algorithms without altering the existing training pipeline. We also introduce a negative sampling strategy to mitigate overestimation errors by reducing rewards for out-of-distribution data, thereby enhancing robustness. This strategy inherently introduces a conservative bias into RL algorithms that employ the learned reward. We evaluate our method on the D4RL benchmark, achieving competitive results compared to offline RL algorithms with access to true rewards and imitation learning (IL) techniques like behavioral cloning. Moreover, our ablation studies on data size and quality reveal the advantages and limitations of SR-Reward as a proxy for true rewards.
@@ 1 Introduction
Imitation learning (IL) from expert demonstrations is one of the most popular avenues for tackling sequential decision-making tasks. There are two categories of methods that make use of expert demonstrations. The first focuses on learning a policy that resembles the expert behavior, e.g. using Behavioral Cloning Pomerleau (1991). Another set of methods, known as inverse reinforcement learning (IRL) (Ng & Russell, 2000), first infer a reward function that explains the expert behavior and then learn a policy derived from that reward function. In popular application domains of IL, such as robotics or medicine, interacting with the environment during training is not always possible due to risk and safety concerns, making it crucial to be able to learn from only the limited, previously collected expert demonstrations.
@@
In this work, we focus on the offline inverse reinforcement learning setting, where the agent neither has access to the reward function nor can query the expert for any feedback. Furthermore, the transition dynamics of the environment are unknown and the agent is provided with limited data in the form of expert demonstrations. Our first contribution, SR-Reward, is a reward function based on Successor Representations (SR), that is learned offline from expert demonstrations. Unlike adversarial schemes popular with IRL methods, our reward function is decoupled from the policy that is being learned. Decoupling the reward from the policy eliminates the instabilities associated with adversarial training and enables the use of a wide range of RL and offline RL algorithms that were previously unusable due to the inaccessibility of the reward function. Moreover, hand-engineering a reward function for complex real-world tasks, such as robot manipulation, is both challenging and error-prone (Wu et al., 2022; Singh et al., 2009). In contrast, demonstrating the desired behavior, while potentially costly, is generally more straightforward. In such scenarios, the ability to learn a dense reward function from demonstrations can be incredibly valuable. Leveraging the SR structure allows SR-Reward to be learned via the Bellman equation which propagates information about future states and actions through temporal difference (TD) learning. Consequently, it can be integrated into existing training pipelines alongside other RL methods based on TD learning with minimal modifications. Simple imitation learning methods like behavioral cloning (BC) directly mimic expert behavior without modeling how actions lead to future states. As a result, BC tends to optimize for short-term objectives and is prone to distribution shift when encountering unseen states. SR-Reward’s integration of SR and TD learning mitigates this issue by enabling a long-term view of the task, making it more resilient to out-of-distribution scenarios.
@@
Function approximation can lead to a significant overestimation of values for out-of-distribution data (Thrun & Schwartz, 1999). Our second contribution is a negative sampling strategy designed to counteract the overestimation error in our reward function for out-of-distribution states and actions. This is accomplished by augmenting the Bellman loss for SR-Reward so that reward estimates for out-of-distribution states and actions decrease based on their distance from expert demonstrations. Incorporating negative sampling not only enhances the robustness of the reward function but also introduces a natural conservatism into the value functions and policies that rely on it.
Our key contributions are:
@@
SR-Reward: a reward function based on Successor Representation (SR), that can be learned solely offline using temporal difference (TD) learning simultaneously with other RL algorithms. We further describe our architecture as well as loss functions needed for training SR-Reward using function approximation.
@@
This paper is organized as follows: section 2 discusses similarities and differences between our proposed reward model and other methods of learning or modifying reward functions. Notation and background information used for developing our method is presented in section 3. Our method, SR-Reward, is presented in detail in section 4, including the motivation, design decisions, and implementation details. Section 4.3 describes our novel negative-sampling method used to increase our reward function’s robustness. Finally, section 5 shows the performance of our method on D4RL environments. We use our reward function in combination with different offline RL algorithms and compare their performance against offline RL with access to the true environment reward. In addition, we compare the performance of our reward in combination with offline RL against IL methods such as BC.
2 Related Work
There is also a series of works that modify the existing environment reward to gain improvements during training. Vieillard et al. (2020) suggests adding of the policy that is being learned to the reward in temporal difference (TD) learning. The authors argue that the logarithm of the policy is a strong learning signal as it is available even in a sparse reward setting and since its value is close to zero for optimal actions under optimal policy this does not conflict with the optimal control objective.
@@ 3 Background
We first introduce the notation and provide a more detailed review of concepts from imitation learning and successor representation.
@@ 3.1 Notation
We consider settings where the environment is represented by a Markov Decision Process (MDP) and is defined as a tuple . and represent the continuous state and continuous action spaces respectively. represents the state transition dynamics, represents the reward function, is the discount factor and represents the starting state distribution. In the offline inverse reinforcement learning setting, we only have access to a limited set of expert demonstrations of the form . In this paper, we are focusing on a minimal setting where neither the transition dynamics nor the reward function are known. The goal is to learn a reward function from expert demonstrations such that its corresponding policy performs similarly to that of the expert.
3.2 Imitation Learning via Distribution Matching
Methods like behavioral cloning (BC), which directly learn a policy mapping states to actions, are straightforward and effective when ample data is available. However, they are prone to distribution shift because they only match the observed action distribution. During inference, as the distribution of encountered states deviates from those seen during training, the accuracy of action predictions diminishes. This leads to accumulating errors that the policy cannot correct.
Distribution matching methods, and related approaches (Ke et al., 2020; Kostrikov et al., 2020; Nachum et al., 2019; Ho & Ermon, 2016; Fu et al., 2018; Ghasemipour et al., 2019), are more robust to distribution shifts since they aim to match both the state and action distributions encountered during training. This helps keep the policy close to the states observed in demonstrations.
where is the indicator function, which equals one if the condition is met and zero otherwise. This is closely related to the state-action distribution . As shown by Puterman (1994), there is a one-to-one correspondence between the state-action distribution and the policy.
Distribution matching methods aim to indirectly learn a policy by minimizing the divergence between and . A common choice is KL-Divergence, and minimizing can be viewed as maximizing the RL objective
where the reward is given by the log ratio of the state-action distributions between the expert policy and the learned policy . Since the state-action distribution is often unavailable, efforts are typically focused on estimating the ratio of the two distributions (Ho & Ermon, 2016; Nachum et al., 2019).
In this paper, we propose a method to directly estimate a proxy for the expert’s state-action distribution from demonstrations and use it as the reward for downstream RL algorithms. We train our reward network using TD learning which can be integrated seamlessly into RL training pipeline and allows for fast inference in continuous high-dimensional spaces.
@@ 3.3 Successor Representations
However, directly estimating SR using these formulations becomes computationally intractable as the number of states and actions increases, or when transitioning from discrete to continuous domains. To address this, previous research (Kulkarni et al., 2016; Machado et al., 2020; Zhang et al., 2017) has extended SR to continuous state and action spaces using Successor Features Representation (SF). SF is expressed in terms of state and action features :
The choice of feature extractor is a design decision that depends on the environment. Most existing work focuses on extracting features only from the state, not the actions. In this scenario, can be represented as , which is a concatenation of state features and actions.
In this work, we adopt this approach and use a feature extractor network to derive features from the state only.
3.3.1 Relationship to State-Action Visitation
SR implicitly captures the state-action visitation. Many density-based IL methods, such as GAIL (Ho & Ermon, 2016), use state-action distribution or occupancy measure for their distribution matching techniques. In Appendix:A, we derive the following close relationship between the occupancy measure and the successor representation
The occupancy measure can be seen as the expectation of successor representations with respect to the probability of all state-action pairs that preceded . We learn successor features representation (SF) from the expert demonstrations using function approximation and use it as a proxy to the occupancy measure of the expert.
4 Technical Approach
@@ 4.3 Negative Sampling
Neural networks tend to overestimate the value of out-of-distribution data points (Thrun & Schwartz, 1999; Ball et al., 2023; Fujimoto et al., 2019; 2018). The overestimation error is especially concerning in our setup because an overestimated value of the reward for unseen states and actions will encourage the value networks and subsequently the policy to diverge from the expert demonstrations. Motivated by the idea of conservative value function via negative sampling Luo et al. (2020), we develop our negative sampling strategy to combat the overestimation error of our SR network. Similar to Luo et al. (2020) we create our negative samples and by adding a small Gaussian noise to states and actions from our expert trajectories. However, instead of subtracting the -norm of the difference vector we decay the values using a Gaussian
with controlling the strength of the decay. Furthermore, we apply negative sampling not to the space of value functions but to the space of rewards. This is possible in our setting where a reward function is estimated and used for learning a policy. Having control over the reward function in this setting provides the opportunity to build conservatism directly into the value functions and subsequently the policy by modifying the reward instead of forcing the value function or the policy to act conservatively Kumar et al. (2020); Fujimoto et al. (2019; 2018).
@@
Figreffig:neg-samp shows the effect of our negative sampling strategy on a toy environment. We train an SR network using ten demonstrations with and without negative sampling and evaluate the rewards over the grid space for each one of the cardinal directions. The mean value plot in Figure 4 shows how negative sampling during training prevents overestimation error for the rewards of the state-action pairs not seen in the demonstrations. Plots for the four main directions show higher reward estimates for the corresponding direction of movement. For example, the Left plot shows the reward for moving left at every grid point and as expected this reward is higher for the upper portion of the trajectories where the expert has moved left.
To further illustrate the effects of negative sampling on preventing over-estimation error, Figure 3 compares the mean return estimated by SR-Reward trained with or without negative sampling. Expert trajectories are corrupted by varying levels of Gaussian noise and their episodic return is estimated using SR-Reward. The model trained with negative sampling produces significantly smaller returns for corrupted trajectories. This effect is shown by an initial drop in the returns when the trajectories are corrupted by Gaussian noise with and a much steeper drop when the noise is increased further. The model trained without negative sampling shows similar return values for expert and corrupted trajectories. The returns for this model show a small drop only after adding a significant amount of noise. Different environments may require different levels of sensitivity to noise and out-of-distribution data. Our negative sampling strategy can be configured for each environment by tuning the level of perturbation noise added to the expert trajectories as well as for the decay rate in Equation 2.
@@ 4.4 Training
We employ several loss functions to train our SR network. As mentioned in Section 3.3, we can estimate the SR using the Bellman equation in a continuous state-action setting. The reward for the Bellman target in Equation 1 is replaced with which is the concatenation of the encoded state and the action. Note that and are calculated without the gradient. We use the -loss to minimize the Bellman error:
To help train the encoder we use an auxiliary prediction task that predicts the next encoded state from the current encoded state and action . We compute the -loss as
We have added an extra loss to penalize the magnitude of the reward for values greater than 1. This loss was found to stabilize the training and create a soft upper bound for the reward. As explained in Section 4.2 we use the -norm of the SR vector as our reward.
Finally, we add a negative sampling loss to improve the robustness of the reward function for out-of-distribution state-action pairs. Similar to Luo et al. (2020) we create negative samples and by perturbing states and actions from the demonstrations with noise. There might be concerns that the negative samples will fall into the same distribution as the demonstrations and so harm the estimation of the SR. However, as discussed by Luo et al. (2020), the demonstrations cover only a small subset of the space, hence the negative samples are with high probability orthogonal to the demonstrations, an effect that increases with the state and action dimensions of the environment. We use isotropic Gaussian noise to create the negative samples. The hyperparameter controls the standard deviation of the Gaussian noise. Intuitively, we want perturbed state-action pairs to have lower reward values proportional to the distance from their counterpart from the dataset. Since SR estimates the visitation count based on , we measure the distance between the negative samples and their original counterparts in the space of features and actions . We calculate the decay factor using an exponential kernel as

### v2

@@ Abstract
In this paper, we propose a novel method for learning reward functions directly from offline demonstrations. Unlike traditional inverse reinforcement learning (IRL), our approach decouples the reward function from the learner’s policy, eliminating the adversarial interaction typically required between the two. This results in a more stable and efficient training process. Our reward module, SR-Reward, leverages successor representation (SR) to encode a state based on expected future states’ visitation under the demonstration policy and transition dynamics. By utilizing the Bellman equation, SR-Reward can be learned concurrently with most reinforcement learning (RL) algorithms without altering the existing training pipeline. We also introduce a negative sampling strategy to mitigate overestimation errors by reducing rewards for out-of-distribution data, thereby enhancing robustness. This strategy introduces an inherent conservative bias into RL algorithms that employ the learned reward, encouraging them to stay close to the demonstrations where the consequences of the actions are better understood. We evaluate our method on D4RL as well as Maniskill Robot Manipulation environments, achieving competitive results compared to offline RL algorithms with access to true rewards and imitation learning (IL) techniques like behavioral cloning.
@@ 1 Introduction
Imitation learning (IL) from expert demonstrations is a widely used approach for tackling sequential decision-making tasks. Methods in this domain generally fall into two categories. The first focuses on directly learning a policy that mimics expert behavior, such as Behavioral Cloning (BC) (Pomerleau, 1991). The second, inverse reinforcement learning (IRL) (Ng & Russell, 2000), infers a reward function that explains an expert’s behavior and simultaneously derives a policy from it. While both approaches have shown promise, they come with notable limitations. BC struggles with distribution shift, failing when encountering states not covered in the demonstrations (Ross et al., 2010). IRL, while more flexible, often inherits the instabilities of adversarial training (Goodfellow et al., 2014) and requires environment interaction to refine the learned reward function — an impractical requirement in domains where exploration is costly or unsafe, such as robotics and healthcare.
Offline reinforcement learning (RL) provides an alternative by learning policies from fixed datasets without interacting with the environment (Lange et al., 2012). However, offline RL typically requires well-defined reward signals, which are often unavailable or difficult to engineer in real-world applications. Without a reward function, the agent has no clear objective and fails to learn a meaningful policy.
To address this challenge, we propose SR-Reward, a reward module that enables the use of offline RL agents in the absence of explicit reward functions. SR-Reward is based on the observation that frequently occurring features in expert demonstrations often capture essential aspects of expertise. By leveraging successor representations (SR)(Dayan, 1993) to estimate feature frequencies, SR-Reward assigns higher rewards to frequently observed state-action pairs. This provides a structured learning signal that helps agents align with expert strategies while retaining the flexibility to adapt to novel situations.
Successor representation provides a way to represent a state based on the frequency of its future visitation. In this work, we leverage SR as the primary mechanism to track how often state-action features appear in the demonstrations. Leveraging the SR structure allows SR-Reward to be learned via the Bellman equation which propagates information about future states and actions through temporal difference (TD) learning. Consequently, it can be integrated into existing training pipelines alongside other RL methods based on TD learning with minimal modifications. Figure 1 illustrates the role of SR-Reward within the training pipeline, highlighting how it integrates into the overall learning process. Unlike adversarial schemes popular with inverse reinforcement learning (IRL) methods (Ng & Russell, 2000; Abbeel & Ng, 2004; Ho & Ermon, 2016), our reward function is decoupled from the policy that is being learned. Decoupling the reward from the policy eliminates the instabilities associated with adversarial training and enables the use of a wide range of RL algorithms that were previously unusable due to the inaccessibility of the reward function (Fujimoto et al., 2019; Garg et al., 2023; Sikchi et al., 2023; Xu et al., 2023).
@@
Hand-engineering a reward function may work for simple scenarios, but it becomes both challenging and error-prone in complex real-world tasks, such as robot manipulation (Singh et al., 2009; Wu et al., 2022). In contrast, demonstrating the desired behavior is generally more straightforward (Wu et al., 2023; Arunachalam et al., 2022; Rakita et al., 2017). It is in such scenarios, that SR-Reward’s ability to learn a dense reward function from demonstrations can be valuable. Another avenue for using the demonstrations is to copy the expert’s actions. Simple imitation learning methods like behavioral cloning (BC) directly mimic expert behavior without modeling how actions lead to future states. Unfortunately, this short-sighted objective is prone to distribution shift when encountering unseen states. SR-Reward enables the use of TD learning methods, which mitigate this issue by gaining a long-term view of the task via bootstrapping, making them more resilient to out-of-distribution scenarios.
@@
We use function approximation to implement the SR mechanism in continuous state and action settings. Function approximation can lead to a significant overestimation of values for out-of-distribution data (Thrun & Schwartz, 1999). This is particularly problematic in our setting as the expert demonstrations by definition cover only a narrow subset of the overall space. We introduce a negative sampling strategy designed to counteract the overestimation error in our reward function for out-of-distribution states and actions. This is accomplished by augmenting the Bellman loss for SR-Reward so that reward estimates for out-of-distribution states and actions decrease based on their distance from expert demonstrations. Incorporating negative sampling not only enhances the robustness of the reward function but also introduces a natural conservatism into the value functions and policies that rely on it. This approach encourages agents to remain near the demonstrated behaviors, where the outcomes of their actions are better understood.
In this work, we focus on the offline inverse reinforcement learning setting, where the agent neither has access to the reward function nor can query the expert for any feedback. Furthermore, the transition dynamics of the environment are unknown and the agent is provided with limited data in the form of expert demonstrations.
In summary our key contributions are:
@@
SR-Reward: a reward function based on Successor Representation (SR), that can be learned solely offline using temporal difference (TD) learning simultaneously with other RL algorithms. We further describe our architecture as well as loss functions needed for training SR-Reward.
@@ 2 Background
We first introduce the notation and provide a more detailed review of concepts from successor representation and imitation learning.
@@ 2.1 Notation
We consider settings where the environment is represented by a Markov Decision Process (MDP) and is defined as a tuple . and represent the continuous state and continuous action spaces respectively. represents the state transition dynamics, represents the reward function, is the discount factor and represents the starting state distribution. In the offline inverse reinforcement learning setting, we only have access to a limited set of expert demonstrations of the form . In this paper, we are focusing on a limited setting where neither the transition dynamics nor the reward function are known. The goal is to learn a reward function from expert demonstrations such that its corresponding policy performs similarly to that of the expert.
@@ 2.2 Successor Representations
However, directly estimating SR using these formulations becomes computationally intractable as the number of states and actions increases, or when transitioning from discrete to continuous domains. To address this, previous research (Kulkarni et al., 2016; Machado et al., 2020; Zhang et al., 2017) has extended SR to continuous state and action spaces using Successor Features Representation (SF). SF is expressed in terms of state and action features vector :
Here is the expected future occupancy of the state-action features for where each element of this vector can be seen as tracking the SR for a single feature of the state-action space. The choice of feature extractor is a design decision that depends on the environment. Most existing work focuses on extracting features only from the state, not the actions. In this scenario, can be represented as , which is a concatenation of state features and actions. In this work, we adopt this approach and use a feature extractor network to derive features from the state only.
2.3 Imitation Learning via Distribution Matching
Methods like behavioral cloning (BC), which directly learn a policy mapping states to actions, are straightforward and effective when ample data is available. However, they are prone to distribution shift because they only match the observed action distribution (Ross et al., 2010). During inference, as the distribution of encountered states deviates from those seen during training, the accuracy of action predictions diminishes. This leads to accumulating errors that the policy cannot correct.
Distribution matching methods, and related approaches (Ho & Ermon, 2016; Fu et al., 2018; Nachum et al., 2019; Ghasemipour et al., 2019; Ke et al., 2020; Kostrikov et al., 2020), are more robust to distribution shifts since they aim to match both the state and action distributions encountered during training. This helps keep the policy close to the states observed in demonstrations.
where is the indicator function, which equals one if the condition is met and zero otherwise. This is closely related to the state-action distribution . Considering that there is a one-to-one correspondence between the state-action distribution and the policy (Puterman, 1994), distribution matching methods aim to indirectly learn a policy by minimizing the divergence between and . A common choice is KL-Divergence, and minimizing can be viewed as maximizing the following RL objective where the reward is given by the log ratio of the state-action distributions between the expert policy and the learned policy (Kostrikov et al., 2020).
Since the state-action distribution is often unavailable, efforts are typically focused on estimating the ratio of the two distributions (Ho & Ermon, 2016; Nachum et al., 2019).
In this paper, we propose a method that estimates SR as a proxy for the expert’s state-action distribution from demonstrations and uses it as a reward for downstream RL algorithms.
2.3.1 Relationship between SR and State-Action Visitation
SR implicitly captures the state-action visitation. Many density-based IL methods, such as GAIL (Ho & Ermon, 2016), use state-action distribution or occupancy measure for their distribution matching techniques. In Appendix:C, we derive the following close relationship between the occupancy measure and the successor representation
The occupancy measure can be seen as the expectation of successor representations with respect to the probability of all state-action pairs that preceded . We learn successor features representation (SF) from the expert demonstrations and use it as a proxy for the expert’s occupancy measure.
3 SR-Reward
@@ 3.3 Negative Sampling
Neural networks tend to overestimate the value of out-of-distribution data points (Thrun & Schwartz, 1999; Fujimoto et al., 2018; 2019; Ball et al., 2023). The overestimation error is especially concerning in our setup because an overestimated value of the reward for unseen states and actions will encourage the value networks and subsequently the policy to diverge from the expert demonstrations. Motivated by the idea of conservative value function via negative sampling Luo et al. (2020), we develop our negative sampling strategy to combat the overestimation error of our SR network. Similar to Luo et al. (2020) we create our negative samples and by adding a small Gaussian noise to states and actions from our expert trajectories. However, instead of subtracting the -norm of the difference vector from the reward estimate of the negative samples, we decay the values using a Gaussian, , with controlling the strength of the decay. Furthermore, we apply negative sampling not to the space of value functions but to the space of rewards. This is possible in our setting where a reward function is estimated and used for learning a policy. Having control over the reward function in this setting provides the opportunity to build conservatism directly into the value functions and subsequently the policy by modifying the reward instead of forcing the value function or the policy to act conservatively Fujimoto et al. (2018; 2019); Kumar et al. (2020).
@@
Figure 5 shows the effect of our negative sampling strategy on a toy environment. We train an SR network using ten demonstrations with and without negative sampling and evaluate the rewards over the grid space for each one of the cardinal directions. The mean value plot in Figure 5 shows how negative sampling during training prevents overestimation error for the rewards of the state-action pairs not seen in the demonstrations. Plots for the four main directions show higher reward estimates for movement in the corresponding direction. For example, the Left plot shows the reward for moving left at every grid point and as expected this reward is higher for the upper portion of the trajectories where the expert has moved left.
To highlight the impact of negative sampling in mitigating overestimation errors, Figure 4 compares the mean trajectory returns of models trained with and without negative sampling. Expert trajectories are corrupted by varying levels of Gaussian noise, and their episodic returns are estimated using SR-Reward. The model trained with negative sampling exhibits significantly lower returns for corrupted trajectories, reflecting its enhanced sensitivity to out-of-distribution data. This effect is evident in the initial drop in returns when Gaussian noise with is applied, followed by a much steeper decline as the noise level increases. In contrast, the model trained without negative sampling produces similar reward estimates for both expert and corrupted trajectories, with returns showing only a modest decline under substantial noise levels. Different environments may demand varying degrees of sensitivity to noise and out-of-distribution data. Our negative sampling strategy offers flexibility, allowing customization for each environment by adjusting the perturbation noise applied to expert trajectories and fine-tuning for the decay rate in Equation 2.
@@ 3.4 Training
We employ several loss functions to train our SR network. As mentioned in Section 2.2, we can estimate the SR using the Bellman equation in a continuous state-action setting. The reward for the Bellman target in Equation 1 is replaced with which is the concatenation of the encoded state and the action. We use the -loss to minimize the Bellman error:
Auxiliary tasks—such as predicting the reward, next state, or next latent state—have been shown to facilitate the learning of useful latent representations for downstream tasks Fujimoto et al. (2024); Ni et al. (2024). To support encoder training, we incorporate an auxiliary prediction task in which the model learns to predict the next encoded state from the current encoded state and action . We compute the -loss as
Ensuring that the reward is bounded is crucial for the stability of reinforcement learning algorithms. Prior works have addressed this in various ways, such as clipping observed rewards or adding a penalty on the magnitude of the estimated reward Mnih et al. (2015); Garg et al. (2021). In our approach, we introduce an additional loss term that penalizes reward magnitudes exceeding 1, effectively encouraging a soft upper bound on the reward. This regularization was found to improve training stability. As described in Section 3.2, we define the reward as the -norm of the SR vector.
Finally, we add a negative sampling loss to improve the robustness of the reward function for out-of-distribution state-action pairs. Similar to Luo et al. (2020) we create negative samples and by perturbing states and actions from the demonstrations with noise. While one might expect negative samples to overlap with the distribution of demonstrations and affect the estimation of the SR, prior work Luo et al. (2020) shows that demonstrations typically cover only a small subset of the state-action space. As a result, negative samples are highly likely to be orthogonal to the demonstrations, an effect that becomes more pronounced as the dimensionality of the environment increases. We use isotropic Gaussian noise to create the negative samples. The hyperparameter controls the standard deviation of the Gaussian noise. Intuitively, we want perturbed state-action pairs to have lower reward values proportional to the distance from their counterpart from the dataset. Since SR estimates the visitation count based on , we measure the distance between the negative samples and their original counterparts in the space of features and actions . We calculate the decay factor using an exponential kernel as